OpenStudy (anonymous):

The formula for the surface area of a cylinder is A=2pr(r+h).Find h if A=80p and r=5. 1 3 5 6 Question 9 1 points Save Solve: |x| = 5 {5} {-5} {-5, 5} {-1, 5} Question 10 1 points Save Solve: | m + 8| = 16 {8} {-24, 8} {8, 16} {-24}

OpenStudy (anonymous):

A = 2pi(r + h) 80 = 2pi(5 + h) 80 = 6.28(5 + h) 12.73 = 5 + h h = 7.73 Answer: I plugged 7.73 back into the equation, and it is correct. Did you type the answers incorrectly?

OpenStudy (anonymous):

{5}

OpenStudy (anonymous):

{8}

OpenStudy (anonymous):

The formula for the surface area of a cylinder is A=2pr(r+h).Find h if A=80p and r=5. 1 3 5 6

OpenStudy (anonymous):

then its 6

OpenStudy (anonymous):

gotb it .. lol one more and im done ...Solve: |2n - 3| = 5 {-1, 4} {1, 4} {-1, 1} {1, -4}

OpenStudy (anonymous):

1,4

OpenStudy (anonymous):

those were all wrong .....

OpenStudy (ajprincess):

2n-3=5 2n=8 n=4 -2n+3=5 2n=-2 n=-1

OpenStudy (ajprincess):

-1,4

OpenStudy (anonymous):